more factoring trinomial
h^2+12h+27 I have (h-9)(h-3)
x^2+14x+33 and I have (x-3)(x-11)

Answer by checkley75(3666)

H^2+12H+27=(H+9)(H+3) NOTE THE + SIGNS. 9H+3H=12H & 9*3=27
--------------------------------------------
H^2+14X+33=(H+11)(H+3) NOTE THE + SIGNS. 11H+3H=14H & 11*3=33